# Ocean-omni: To Understand the World with Omni-modality

### v3

@@ 4.5.3 Audio-Language Branch
For analysis, we measure the average WER on all our ASR benchmarks across Fleurs, WenetSpeech and KeSpeech by training a 1.5B audio-language model with three different down-sampling rates 2, 4, 8. To simulate the actual training of the audio branch in Ocean-omni , we only train our audio encoder and projector while keeping the LLM frozen. This setup is consistent with the configuration described in Section 3.2.3.
@@ 4.5.4 Multimodal Supervised Fine-Tuning
Table 11 and Table 12 compare the performance of Ocean-omni on various image and video benchmarks with and without multimodal supervised fine-tuning (SFT). The results indicate that the model exhibits superior overall performance after undergoing multimodal SFT compared to the version that only undergoes instruction fine-tuning (IFT). This improvement can be attributed to the use of high-quality, diverse instructions and our SFT data construction method, which avoid compromising the base model’s capabilities. (See Section 3.3 for more details.)
Table 11: Performances of Ocean-omni on image tasks before and after the supervised fine-tuning stage. Generally, the model’s performance has improved across most image benchmarks.
@@ 5 Conclusion
In this work, we have open-sourced Ocean-omni as a step toward developing a truly omni-modal LLM that encompasses all human senses. With omni-modal pretraining and fine-tuning using high-quality omni-modal data, this version of Ocean-omni has achieved leading levels in integrating comprehension across video, image, text, and audio. Despite its promising performance, there remains significant room for improvement in the foundational capabilities across each individual modality. This include (1) enhancing text extraction capabilities, (2) supporting longer video understanding, (3) developing an end-to-end TTS system integrated with LLMs, and (4) improving the ability to comprehend not only human voices but also natural environmental sounds, such as flowing water, bird calls, and collision noises, among others.

### v4

@@ 4.5.3 Audio-Language Branch
For analysis, we measure the average WER on all our ASR benchmarks across Fleurs, WenetSpeech and KeSpeech by training a 1.5B audio-language model with three different down-sampling rates 2, 4, 8. To simulate the actual training of the audio branch in Baichuan-omni , we only train our audio encoder and projector while keeping the LLM frozen. This setup is consistent with the configuration described in Section 3.2.3.
@@ 4.5.4 Multimodal Supervised Fine-Tuning
Table 11 and Table 12 compare the performance of Baichuan-omni on various image and video benchmarks with and without multimodal supervised fine-tuning (SFT). The results indicate that the model exhibits superior overall performance after undergoing multimodal SFT compared to the version that only undergoes instruction fine-tuning (IFT). This improvement can be attributed to the use of high-quality, diverse instructions and our SFT data construction method, which avoid compromising the base model’s capabilities. (See Section 3.3 for more details.)
Table 11: Performances of Baichuan-omni on image tasks before and after the supervised fine-tuning stage. Generally, the model’s performance has improved across most image benchmarks.
@@ 5 Conclusion
In this work, we have open-sourced Baichuan-omni as a step toward developing a truly omni-modal LLM that encompasses all human senses. With omni-modal pretraining and fine-tuning using high-quality omni-modal data, this version of Baichuan-omni has achieved leading levels in integrating comprehension across video, image, text, and audio. Despite its promising performance, there remains significant room for improvement in the foundational capabilities across each individual modality. This include (1) enhancing text extraction capabilities, (2) supporting longer video understanding, (3) developing an end-to-end TTS system integrated with LLMs, and (4) improving the ability to comprehend not only human voices but also natural environmental sounds, such as flowing water, bird calls, and collision noises, among others.